Two sides of a triangle have the same length. The third side measures 3 m less than twice the common length. The perimeter of the triangle is 25m. What are the lengths of the three sides?
Answer by stanbon(75887)

Two sides of a triangle have the same length. The third side measures 3 m less than twice the common length. The perimeter of the triangle is 25m. What are the lengths of the three sides?
----------
Let the equal sides have length "x".
The 3rd side has meansure "2x-3".
----
Equation:
x + x + 2x-3 = 25
4x = 28
x = 7m
2x-3 = 2*7-3 = 11m
-------------------------
cheers,
Stan H.